# VL-Nav: Real-time Vision-Language Navigation with Spatial Reasoning

### v3

@@ III-D CVL Scoring Policy
As shown in Fig. 3, a vision-language (VL) score $S_{\mathrm{VL}}(\mathbf{g})$ translates pixel-wise vision-language features into a Gaussian-mixture distribution over the robot’s horizontal field of view (FOV). Suppose the open-vocabulary detection model identifies $K$ likely directions or detection regions, each parameterized by $\bigl{(}\mu_{k},\sigma_{k},\alpha_{k}\bigr{)}$. Here, $\mu_{k}$ indicates the mean offset angle within the FOV, $\sigma_{k}$ encodes the detection’s angular uncertainty (in our implementation, a fixed constant of $0.1$ for all detections), and $\alpha_{k}$ is a confidence-based weight that captures how important that region is. By accumulating each region’s contribution in a mixture model, $S_{\mathrm{VL}}(\mathbf{g})$ naturally biases the robot toward goals in the direction of high-confidence detections that match the language prompt.
From a human perspective, this process resembles observing different areas that related to the target instance with varying levels of certainty. For instance, if someone is looking for a “red chair” in a large room, they might spot multiple red silhouettes at the right side of their vision, plus a clearer view of something red near the center. Although peripheral glimpses (“out of the corner of the eye”) often carry lower confidence, they are not ignored. Instead, we naturally merge these partial observations into a mental sense of “the chair is probably in one of these directions.” The system follows a similar approach, assigning each detection region a Gaussian shape around $\mu_{k}$, scaled by $\sigma_{k}$ and weighted by $\alpha_{k}$. Inspired by VLFM [39], we then multiply by $C(\Delta\theta)$, a “view confidence” term that downweights detections with large angular offsets from the central field of view. The VL score is computed as:
@@
(1) Distance Weighting. We define
@@
[…]
@@
[…]
@@
$S_{\mathrm{CVL}}(\mathbf{g})\;=\;\Bigl{(}w_{\mathrm{dist}}\;S_{\mathrm{dist}}(\mathbf{g})\;+\;w_{\mathrm{VL}}\;S_{\mathrm{VL}}(\mathbf{g})\;\cdot\;S_{\mathrm{unknown}}(\mathbf{g})\Bigr{)},$
where $w_{\mathrm{VL}}$ and $w_{\mathrm{dist}}$ are scalar weights. Note that the VL score is assigned when a point is initially detected, whereas the curiosity terms (distance and unknown-area weighting) are evaluated at goal selection time—when the current distance and unknown ratio can be calculated. For instance-based goals, these curiosity cues are omitted because their primary purpose is verification rather than exploration.
III-E Select & Publish Goal Point
After evaluating CVL scores for all instance-based and frontier goals, the system picks the highest-scoring candidate. Algorithm 2 sketches the selection process, prioritizing instances first, then frontiers.
Algorithm 2 Select and Publish Goal Point
0:  Instance points $\{\mathbf{o}_{1},\dots\}$ with $VL\_score$, frontier points $\{\mathbf{f}_{1},\dots\}$ with $CVL\_score$, robot position $\mathbf{r}$, and threshold $\delta_{\mathrm{reached}}$ for discarding points too close to $\mathbf{r}$.
1:  $\textit{bestScore}\leftarrow-\infty$
[…]
3:  for all instances $\mathbf{o}_{i}$ do
4:     if $\|\mathbf{o}_{i}-\mathbf{r}\|>\delta_{\mathrm{reached}}$ then
[…]
6:        if $s>\textit{bestScore}$ then
7:           $\textit{bestScore}\leftarrow s$
8:           $\textit{bestGoal}\leftarrow\mathbf{o}_{i}$
9:        end if
10:     end if
11:  end for
[…]
13:     for all frontiers $\mathbf{f}_{j}$ do
14:        if $\|\mathbf{f}_{j}-\mathbf{r}\|>\delta_{\mathrm{reached}}$ then
[…]
16:           if $s>\textit{bestScore}$ then
17:              $\textit{bestScore}\leftarrow s$
18:              $\textit{bestGoal}\leftarrow\mathbf{f}_{j}$
19:           end if
20:        end if
21:     end for
22:  end if
[…]
24:     Publish bestGoal to /selected_goal_point
25:  end if
26:  return
If no candidate exceeds the distance threshold $\delta_{\mathrm{reached}}$, the robot withholds any new goal. Once a goal is published, a global planner updates obstacle polygons and replans a safe route. A local planner refines each global waypoint into short-horizon commands for real-time obstacle avoidance, supporting incremental exploration in unknown environments.
III-F Path Planning with FAR Planner
Once an CVL goal is selected, we hand off point-goal path planning to the FAR Planner [36], which represents obstacles polygonally and updates a visibility graph in real time. This enables efficient re-planning in partially unknown environments, often outperforming search- or sampling-based methods. A local planner refines FAR Planner’s waypoints into short-horizon velocity commands, ensuring swift reactions to new obstacles and seamless integration with VL-Nav’s zero-shot vision-language objectives.
@@ IV Experiments
Figure 4: Four different real-world experiment environments.
(a) Office
(b) Apartment
(c) Outdoor
(d) Hallway
Figure 5: Top-down view of the trajectories comparison on the value maps with the detection results across the four different environments.
TABLE I: Vision-language navigation performance in 4 unseen environments (SR and SPL).
Method SR (%) SPL
@@
IV-A Experimental Setting
We evaluate our approach in real-robot experiments against five methods: (1) classical frontier-based exploration, (2) VLFM [39], (3) VLNav without instance-based target points, (4) VLNav without curiosity terms, and (5) the full VLNav configuration. Because the original VLFM relies on BLIP-2 [15], which is too computationally heavy for real-time edge deployment, we use the YOLO-World [6] model instead to generate per-observation similarity scores for VLFM. Each method is tested under the same conditions to ensure a fair comparison of performance.
@@ IV-A Experimental Setting
We consider four distinct environments (shown in Fig. 4), each with a specific combination of semantic complexity (High, Medium, or Low) and size (Big, Mid, or Small). Concretely, we use a Hallway (Medium & Big), an Office (High & Mid), an Outdoor area (Low & Big), and an Apartment (High & Small). In each environment, we evaluate five methods using three language prompts, yielding a diverse range of spatial layouts and semantic challenges. This setup provides a rigorous assessment of each method’s adaptability.
Language-Described Instance:
@@
We define nine distinct, uncommon human-described instances to serve as target objects or persons during navigation. Examples include phrases such as “tall white trash bin,” “there seems to be a man in white,” “find a man in gray,” “there seems to be a black chair,” “tall white board,” and “there seems to be a fold chair.” The variety in these descriptions ensures that the robot must rely on vision-language understanding to accurately locate these targets.
Robots and Sensor Setup: All experiments are conducted using a four-wheel Rover equipped with a Livox Mid-360 LiDAR. The LiDAR is tilted by approximately 23 degrees to the front to achieve a ±$\pm 30$ degrees vertical FOV coverage closely aligned with the forward camera’s view. An Intel RealSense D455 RGB-D camera, tilted upward by 7 degrees to detect taller objects, provides visual observation, though its depth data are not used for positioning or mapping. LiDAR measurements are a primary source of mapping and localization due to their higher accuracy. The whole VL-Nav system runs on an NVIDIA Jetson Orin NX on-board computer.
IV-B Main Results
We validate the proposed VL-Nav system in real-robot experiments across four distinct environments (Hallway, Office, Apartment, and Outdoor), each featuring different semantic levels and sizes. Building on the motivation articulated in Section I, we focus on evaluating VL-Nav’s ability to (1) interpret fine-grained vision-language features and conduct robust VLN, (2) explore efficiently in unfamiliar spaces across various environments, and (3) run in real-time on resource-constrained platforms. Fig. 5 presents a top-down comparison of trajectories and detection results on the value map.
Figure 6: Plots of performance in different environments sizes and semantic comlexities.
Overall Performance:
As reported in Table I, our full VL-Nav consistently obtains the highest Success Rate (SR) and Success weighted by Path Length (SPL) across all four environments. In particular, VL-Nav outperforms classical exploration by a large margin, confirming the advantage of integrating CVL spatial reasoning with partial frontier-based search rather than relying solely on geometric exploration.
Effect of Instance-Based Target Points (IBTP):
We note a marked improvement when enabling IBTP: the variant without IBTP lags behind, particularly in complex domains like the Apartment and Office. As discussed in Section III, IBTP allows VL-Nav to pursue and verify tentative detections with confidence above a threshold, mirroring human search behavior. This pragmatic mechanism prevents ignoring possible matches to the target description and reduces overall travel distance to confirm or discard candidate objects.
Curiosity Contributions:
The curiosity Score is also significant to VL-Nav’s performance. It merges two key components:
  • Distance Weighting: Preventing easily select very far way goals to reduce travel time and energy consumption which is extremely important for the efficiency (metrics SPL) in the large-size environments.
  • Unknown-Area Weighting: Rewards navigation toward regions that yield more information.
Our ablations reveal that removing the distance-scoring element (VL-Nav w/o curiosity) degrades both SR and SPL, particularly in the more cluttered environments. Meanwhile, dropping the instance-based target points (IBTP) similarly lowers performance, reflecting how each piece of CVL addresses a complementary aspect of semantic navigation.
Comparison to VLFM:
Although the VLFM approach [39] harnesses vision-language similarity value, it lacks the pixel-wise vision-language features, instance-based target points verification mechanism, and CVL-based spatial reasoning. Consequently, VL-Nav surpasses VLFM in both SR and SPL by effectively combining the pixel-wise vision language features and the curiosity cues via the CVL spatial reasoning. These gains are especially pronounced in semantic complex (Apartment) and open-area (Outdoor) environments, underscoring how our CVL spatial reasoning enhance vision-language navigation in complex settings and scenarios.
Summary of Findings:
In conclusion, the experimental results confirm that VL-Nav delivers superior vision-language navigation across diverse, unseen real-world environments. By fusing frontier-based target points detection, instance-based target points, and the CVL spatial reasoning for goal selection, VL-Nav balances semantic awareness and exploration efficiency. The system’s robust performance, even in large or cluttered domains, highlights its potential as a practical solution for zero-shot vision-language navigation on low-power robots.
V Discussion
V-A Sizes of Environments vs. SR and SPL
Our experiments grouped environments by size into three categories: Small (Apartment), Mid (Office), and Big (averaging Hallway and Outdoor). As shown in Table I and the line plots in Fig. 6, classical Frontier Exploration suffers a marked performance decline in larger spaces, frequently exceeding the path length limit. Its success rate drops from about 55.6% in the Small environment to only 36.7% in the Big, and its SPL falls from 0.564 to 0.214. This illustrates how naive 360° frontier-based approaches grow increasingly inefficient in larger, more open environments.
By contrast, VL-Nav remains robust across all sizes. Although there is a slight decrease when moving from Mid (91.7% SR, 0.812 SPL) to Big (82.3% SR, 0.655 SPL), its overall performance remains substantially higher than the baselines. A similar trend appears for the ablated VL-Nav variants (w/o IBTP, w/o curiosity), which still surpass Frontier Exploration but fall behind the full VL-Nav system.
These findings underscore two key factors. First, CVL spatial reasoning allows our method to leverage fine-grained vision-language cues (e.g., the “gray cloth” in Fig. 1) while also prioritizing areas with substantial unknown space, thus avoiding the extreme inefficiencies of frontier-based exploration faces in large, open environments. Second, instance-based target points (IBTP) enable more targeted navigation, sustaining high success rates even as map size grows.
V-B Semantic Complexities of Environments vs. SR and SPL
We also vary the environments by semantic complexity, classifying them as Low (Outdoor), Medium (Hallway), and High (averaging Office and Apartment). An interesting trend emerges: all methods show better performance in more semantically rich environments. The likely explanation is that structured indoor spaces (e.g., with walls or furniture) provide stronger cues for the detection and segmentation models used by all five methods, leading to more accurate navigation.
These results underscore that VL-Nav gains more pronounced benefits as the environment offers more semantic context. Rather than being overwhelmed by clutter or complexity, the system capitalizes on language-guided instance-based detections and curiosity-based exploration . The fusion of the CVL spatial reasoning and instance-based target points provides a scalable strategy to locate human-specified targets, even in highly complex indoor domains.
V-C Dynamic Environment Handling
Handling dynamic environments is crucial for robust navigation in real-world scenarios, where changes such as moving obstacles and varying terrain can impact robot performance. VL-Nav incorporates several strategies to dynamically adapt to environmental changes, ensuring reliability and efficiency.
To handle the dynamic environment, we maintain a dynamic occupancy grid map that integrates dynamic terrain analysis and path planning within the FAR planner [36]. This allows the system to continuously update its understanding of the environment and adjust navigation strategies accordingly. The dynamic grid map ensures that real-time changes, such as shifting obstacles and terrain conditions, are accurately reflected, providing a more robust foundation for dynamic vision-language navigation in the social environment.
V-D Real-Robot Efficiency Optimization
Although VL-Nav is designed as a general navigation framework, we introduce several optimizations to ensure real-world feasibility and high performance on mobile robots.
Visual-Language Module Variants:
The first optimization involves selecting an efficient model variant for the visual-language (or object-detection) pipeline. In our tests, we ran multiple versions of YOLO-World [6] detectors with different input dimensions and TensorRT precision modes. For instance, one configuration used a $256\times 320$ input size in a standard GPU runtime, another used TensorRT (FP32) at the same resolution, and a third tested TensorRT (FP16) at $480\times 640$. Empirically, the original $256\times 320$ configuration (run without TensorRT) achieved about $23\%$ higher success rates than higher-resolution FP16 modes. This suggests that moderate input resolutions can strike a more favorable balance between detection accuracy and real-time performance, which is crucial for robust and efficient continuous on-robot inference.
Rolling Occupancy Grid vs. Fixed Large Grid:
A second optimization lies in how the occupancy grid is maintained. Rather than keeping a large, fixed-size global grid (common in many simulation environments), we implement a rolling grid on the real robot, dynamically shifting or expanding it only as new sensor data demand. This rolling approach keeps memory usage proportional to the active exploration area, lowering both storage overhead and BFS/cluster computations for partial frontier detection. Moreover, since real-world robots typically explore incrementally, discarding far-off map regions has minimal impact on decision-making.
Additional System-Level Improvements:
Finally, we incorporate various practical measures to ensure robust navigation under real-world constraints. For instance, the robot’s local planner and velocity commands are monitored to detect “stuck” situations; if the robot remains motionless beyond a time threshold, the system discards its current goal and reselects a new one. Similarly, sensor frequency and map update rates are tuned so that occupancy-grid expansions do not saturate available CPU/GPU resources. These measures, alongside the rolling-grid and efficient detection models, help keep VL-Nav lightweight and reactive.
VI Limitation and Conclusion
VI-A Limitations
Despite the promising results demonstrated by VL-Nav, several limitations remain. One major challenge is the system’s ability to handle complex language descriptions. It struggles with spatial descriptions that contain hidden object references and objects described with specific textual annotations, which can affect navigation accuracy. Another limitation lies in the reliance on manually defined thresholds for various navigation conditions like the lightning condition. These thresholds may not generalize well across different environments and scenarios, requiring further investigation into adaptive or learning-based approaches to threshold tuning.
VI-B Conclusion
In this paper, we introduced VL-Nav, a vision-language navigation framework that operates efficiently in real time on resource-constrained platforms. By combining pixel-wise vision-language features with a curiosity-based exploration strategy, our CVL spatial reasoning method demonstrates robust performance in a variety of indoor and outdoor settings. Empirically, VL-Nav not only achieves real-time navigation at 30 Hz on a Jetson Orin NX but also surpasses existing approaches by 44.15%, achieving an overall success rate of 86.3%. The key to these gains lies in effectively leveraging semantic cues from vision-language embeddings to prioritize frontiers and potential object instances, thereby aligning navigation decisions more closely with human-like reasoning.
Future work will explore extending VL-Nav to handle more complex instructions that involve multi-step tasks, temporal reasoning (e.g., tracking a moving target), and dynamic environments. Additionally, further integration with large language models could enable more nuanced command parsing and open-vocabulary object detection for even broader applicability. We believe these directions will help advance the capability of VLN systems and bring us closer to deploying versatile, robust robotic assistants in real-world settings.

### v4

@@ III-B3 NeSy Scoring Policy
As shown in Fig. 4, a vision-language (VL) score $S_{\mathrm{VL}}(\mathbf{g})$ translates fine-grained vision-language features from the YOLO-World [6] and FastSAM [40] models into a Gaussian-mixture distribution over the robot’s horizontal field of view (FoV). Suppose the open-vocabulary detection model identifies $K$ likely directions or detection regions, each parameterized by $\bigl(\mu_{k},\sigma_{k},\alpha_{k}\bigr)$. Here, $\mu_{k}$ indicates the mean offset angle within the FoV, $\sigma_{k}$ encodes the detection’s angular uncertainty (in our implementation, a fixed constant of $0.1$ for all detections), and $\alpha_{k}$ is a confidence-based weight that captures how important that region is. By accumulating each region’s contribution in a mixture model, $S_{\mathrm{VL}}(\mathbf{g})$ naturally biases the robot toward goals in the direction of high-confidence detections that match the subtask prompt.
From a human perspective, this process resembles observing different areas that related to the target instance with varying levels of certainty. For instance, if someone is looking for a “red chair” in a large room, they might spot multiple red silhouettes at the right side of their vision, plus a clearer view of something red near the center. Although peripheral glimpses (“out of the corner of the eye”) often carry lower confidence, they are not ignored. Instead, we naturally merge these partial observations into a mental sense of “the chair is probably in one of these directions.” The system follows a similar approach, assigning each detection region a Gaussian shape around $\mu_{k}$, scaled by $\sigma_{k}$ and weighted by $\alpha_{k}$. Inspired by VLFM [36], we then multiply by $C(\Delta\theta)$, a “view confidence” term that downweights detections with large angular offsets from the central field of view. The VL score is computed as following equations:
@@
Finally, to ensure the metric remains bounded, the accumulated score $S_{\mathrm{VL}}(\mathbf{g})$ is clipped to the range $[0,1]$.
@@
(1) Distance Weighting. We define:
@@
$\mathrm{ratio}(\mathbf{g})\;=\;\frac{\#(\text{unknown cells})}{\#(\text{total visited cells})},$
@@
$S_{\mathrm{unknown}}(\mathbf{g})\;=\;1\;-\;\exp\!\Bigl(-\,k\,\mathrm{ratio}(\mathbf{g})\Bigr),$
@@
$S_{\mathrm{NeSy}}(\mathbf{g})\;=\;\Bigl(w_{\mathrm{dist}}\;S_{\mathrm{dist}}(\mathbf{g})\;+\;w_{\mathrm{VL}}\;S_{\mathrm{VL}}(\mathbf{g})\;\cdot\;S_{\mathrm{unknown}}(\mathbf{g})\Bigr),$
where $w_{\mathrm{VL}}$ and $w_{\mathrm{dist}}$ are scalar weights. Note that the VL score is assigned when a point is initially detected, whereas the curiosity terms (distance and unknown-area weighting) are evaluated at goal selection time, when the current distance and unknown ratio can be calculated. For instance-based goals, these curiosity cues are omitted because their primary purpose is verification rather than exploration.
III-C Goal Selection & Path Planning
After computing scores, the system selects the optimal goal point by prioritizing high-confidence semantic matches over general exploration. Specifically, we iterate through all available instance-based target points that exceed a minimum distance threshold $\delta_{\mathrm{reached}}$. If any valid instance targets exist, the one with the highest $S_{\mathrm{VL}}$ is selected immediately to drive verification. If no valid instance targets are available (or all are too close), the system falls back to frontier-based target points. The frontier with the highest combined $S_{\mathrm{NeSy}}$ score is chosen to maximize information gain. If no candidate exceeds $\delta_{\mathrm{reached}}$, the robot holds its position. Once a goal is selected, we employ the FAR Planner [34] to generate a collision-free path, guiding the robot to the target while dynamically avoiding real-time obstacles.
@@ IV Experiments
IV-A Simulation Experiments: DARPA TIAMAT Challenge
We first validate VL-Nav in the high-fidelity simulation environment of the DARPA TIAMAT Challenge (Phase 1). This benchmark is designed to test robotic reasoning under highly abstract and complex human instructions.
Baselines:
We compare VL-Nav against four well-known baselines and two ablation variants: closest frontier exploration, VLFM [36], SG-Nav [35], ApexNav [39], VL-Nav w/o IBTP (ablation), and VL-Nav w/o Curiosity (ablation).
Environments & Simulators:
The evaluation covers four distinct scenarios simulated across two platforms:
  • HabitatSim (Indoor): The Apartment 1 and Apartment 2 environments are rendered in HabitatSim [29] to evaluate navigation in daily residential settings.
  • IsaacSim (Outdoor): The Camping Site and Factory environments are simulated in IsaacSim [23] to leverage its advanced physics engine for unstructured settings.
Task Settings:
Except for interaction, each environment consists of 8 abstract navigation tasks requiring multi-step reasoning. To ensure reliability, we conduct 3 trials for all tasks in each environment. Crucially, some tasks require the robot to locate multiple target objects distributed across widely separated positions within a large-scale map. Instructions involve implicit semantics (e.g., “it is raining” $\rightarrow$ find umbrella), hierarchical searching (e.g., “check the garage for tools”), and abstract social concepts (e.g., “prepare for a black-tie party” $\rightarrow$ find suit, shoes, and tie).
Robot and Sensor Setup: The DARPA TIAMAT tasks require a Boston Dynamics Spot robot, equipped with five RGB-D cameras positioned to provide a comprehensive surround view for visual reasoning.
TABLE I: Simulation experiment results on the navigation tasks of DARPA TIAMAT challenge phase 1 in 4 different scenarios with metrics Success Rate (SR) and Max Time Usage Ratio (MTUR).
Method SR % ($\uparrow$) MTUR ($\downarrow$)
Apartment 1 Apartment 2 Camping Site Factory Apartment 1 Apartment 2 Camping Site Factory
Frontier Exploration 8.3 8.3 0.0 0.0 0.958 0.884 1.000 1.000
VLFM [36] 8.3 8.3 4.2 8.3 0.931 0.864 0.953 0.859
SG-Nav [35] 0.0 4.2 0.0 8.3 1.000 0.973 1.000 0.901
ApexNav [39] 25.0 25.0 20.8 12.5 0.817 0.795 0.828 0.861
VL-Nav w/o IBTP 70.8 62.5 62.5 58.3 0.680 0.724 0.731 0.762
VL-Nav w/o Curiosity 79.1 75.0 58.3 66.7 0.612 0.635 0.793 0.735
VL-Nav 87.5 79.2 75.0 75.0 0.562 0.591 0.647 0.679
Figure 5: Qualitative results in the DARPA TIAMAT Indoor-2 environment. We demonstrate the system’s capability to resolve three distinct types of complex instructions: (1) implicit semantic inference (weather $\rightarrow$ rain jacket); (2) hierarchical task decomposition (garage $\rightarrow$ tools); and (3) abstract concept grounding (black-tie party $\rightarrow$ formal suit).
Figure 6: Qualitative results in the DARPA TIAMAT outdoor factory environment. We demonstrate robust navigation under unstructured conditions across three scenarios: (1) single-object retrieval; (2) implicit inference (measure the perimeter of a pipe $\rightarrow$ measuring tape); and (3) a challenging multi-objective task where the robot must locate three distinct targets scattered across different regions of the large-scale map. The colored point clouds in the insets visualize the value map guiding the exploration.
Figure 7: Top: Four simulation environments in the DARTPA TIAMAT Challenge. Bottom: Four real-world experiment environments.
TABLE II: Vision-language navigation performance in 4 unseen environments (SR and SPL).
Method SR % ($\uparrow$) SPL ($\uparrow$)
@@ IV-A Simulation Experiments: DARPA TIAMAT Challenge
IV-B Real-World Experiments
Baselines:
Since SG-Nav [35] and ApexNav [39] cannot be deployed on the edge device, we evaluate our approach in real-robot experiments against five methods: (1) closet frontier-based exploration, (2) VLFM [36], (3) VL-Nav without instance-based target points, (4) VL-Nav without curiosity terms, and (5) the full VL-Nav configuration. Because the original VLFM relies on BLIP-2 [17], which is too computationally heavy for real-time edge deployment, we use the same YOLO-World [6] and FastSAM [40] models instead to generate per-observation similarity scores for VLFM. Each method is tested under the same conditions to ensure a fair comparison of performance.
@@ IV-B Real-World Experiments
We consider four distinct environments (shown in Fig. 7), each with a specific combination of semantic complexity (High, Medium, or Low) and size (Big, Mid, or Small). Concretely, we use a Hallway (Medium & Big), an Office (High & Mid), an Outdoor area (Low & Big), and an Apartment (High & Small).
Task Settings:
@@
We define nine distinct, uncommon human-described instances to serve as target objects or persons during navigation (e.g., “tall white trash bin,” “there seems to be a man in white”). This results in 9 navigation tasks in total, and we conduct 3 to 5 trials for each task to robustly evaluate performance. This ensures that the robot must rely on NeSy reasoning to accurately locate these targets.
Real Robot Implementation:
Experiments are conducted on a four-wheeled Rover, with a Unitree Go2 quadruped used for the multi-floor demonstration. Both robots are equipped with a Livox Mid-360 LiDAR (tilted $23^{\circ}$) and an Intel RealSense D455 camera (tilted $7^{\circ}$). The Go2 is additionally equipped with a Unitree D1 arm and a RealSense D435 camera to enable object grasping. Computation is distributed: the real-time navigation stack runs onboard on an NVIDIA Jetson Orin NX, while the NeSy Task Planner runs on a remote laptop with an RTX 4090 GPU to host the Qwen3-VL-8B model [1]. This separation is effective as the task planner operates asynchronously, updating only when a new task is received or a subtask is completed.
IV-C Main Results
We validate the proposed VL-Nav system through rigorous testing in both the high-fidelity DARPA TIAMAT simulation and diverse real-world deployments. Our evaluation employs distinct metrics suited for each domain: for simulation, we report Success Rate (SR) and Max Time Usage Ratio (MTUR), defined as the average execution time divided by the maximum time limit per task, to quantify time efficiency and completion reliability. For real-world experiments, we adopt SR and Success weighted by Path Length (SPL) following the VLFM convention [36] to evaluate efficiency.
Simulation Results (DARPA TIAMAT)
As reported in Table I, VL-Nav achieves dominant performance with SRs ranging from 75.0% to 87.5%, whereas baselines struggle.
  • Baseline Failures and High MTUR: Standard approaches fail to resolve the abstract logic and multi-target sequencing required by the benchmark. Notably, SG-Nav [35] exhibits an MTUR near 1.0 across environments, frequently timing out. This is primarily because its reliance on LLM-based Chain-of-Thought reasoning at every navigation step incurs prohibitive latency [35]. Similarly, ApexNav [39] suffers from low success (SR $\approx$ 12.5%–25.0%) as it lacks the consistent symbolic memory to efficiently manage long-horizon and multiple goals located at different positions.
  • Qualitative Reasoning: The system’s ability to ground abstract concepts is visualized in Figures 5 and 6. In indoor settings (Fig. 5), VL-Nav successfully infers implicit semantics (e.g., “weather” $\rightarrow$ “rain jacket”) and performs hierarchical decomposition. In unstructured outdoor settings (Fig. 6), the projected value maps demonstrate how the system targets large landmarks to guide the search for smaller, distributed objects (laptop, outfit, truck) without getting lost in open space.
Real-World Results
Physical deployments (Table II) confirm the system’s robustness and efficiency.
  • Efficiency (SPL): VL-Nav achieves significantly higher SPL scores (e.g., 0.812 in Office) compared to Frontier Exploration (0.317) and VLFM (0.556). This metric confirms that VL-Nav does not merely find targets by chance but navigates via efficient, non-redundant paths, verifying that the neuro-symbolic planning translates effectively to real-world dynamics.
  • Sim-to-Real Transfer: The consistent performance gap between VL-Nav and baselines across both simulation and real-world experiments highlights the system’s strong generalization capabilities.
Ablation Analysis
  • Impact of IBTP (Verification): The w/o IBTP variant suffers most in semantically complex and cluttered environments, such as the Apartment (Sim: 87.5% $\rightarrow$ 70.8%; Real: 88.9% $\rightarrow$ 70.2%). Without IBTP, the robot lacks the “short-cut” mechanism to verify tentative detections, leading to missed targets and lower SPL.
  • Impact of Curiosity (Exploration): The w/o Curiosity variant degrades primarily in large-scale, open environments, such as the Factory and Outdoor scenes (Sim Factory: 75.0% $\rightarrow$ 66.7%; Real Outdoor: 77.8% $\rightarrow$ 55.6%). This confirms that distance and unknown-area weighting are critical for preventing local minima and aimless wandering in expansive maps.
Summary
The combined results demonstrate that while baselines like SG-Nav and ApexNav stumble on the computational and logical demands of abstract multi-target tasks, VL-Nav’s neuro-symbolic architecture provides the necessary reasoning depth and execution efficiency to succeed in both high-fidelity simulation and real-world environments.
V Conclusions and Future Work
In this paper, we presented VL-Nav, a neuro-symbolic framework that bridges the gap between abstract human instructions and robotic execution. By intertwining a NeSy Task Planner with a NeSy Exploration System, our approach effectively addresses the logical and spatial reasoning challenges inherent in large-scale vision-language navigation. Extensive validation across the DARPA TIAMAT challenge and diverse real-world experiments demonstrates the system’s robustness, achieving overall success rates of 79.2% and 86.3%, respectively. These results confirm VL-Nav’s capability to complete complex tasks in large-scale unseen environments. Future work will focus on extending the symbolic memory to support temporal reasoning, enabling the robot to track moving targets and adapt to dynamic environmental changes. Additionally, we aim to incorporate lifelong learning capabilities, allowing the system to adapt to evolving environments over long-term deployments.